The product of two consecutive odd integers is 63. Find the integers.
Answer by checkley77(12569)

x(x+2)=63
x^2+2x-63=0
(x+9)(x-7)=0
x+9=0
x=-9 answer.
-9+2=-7 answer.
proof
-9*-7=63
63=63